OpenStudy (anonymous):

Show that; y'' = -2x/y^5 , when x^3 - y^3 =1.

OpenStudy (amistre64):

define y' for me

OpenStudy (amistre64):

or just take a first derivative of the given equation, and then a second, defining y' will help simplify the second derivative overall

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

x^3 - y^3 =1 3x^2 - 3y^2 y' = 0 6x - 6y y' - 3y^2 y'' = 0 the rest is simple algebra

OpenStudy (anonymous):

That was close, I also had the same answer like yours. But that does not complete the proof.

OpenStudy (amistre64):

im not here to complete your proof, i am here to guide you along the way.

OpenStudy (amistre64):

the last setup has all the needed information for you to work with, just algebra your way thru it and simplify the results